OpenStudy (anonymous):

Please Answer this my friend needs a help, Twice the smaller number of nine less than the larger number if the difference of the two numbers is 15 find the numbers.

OpenStudy (anonymous):

let smaller number = x and larger = y then do u mean 'smaller number IS nine less than larger number'?

OpenStudy (sriram):

X-Y=15 and x-2y=9 hence y=6 (I - II) and x=21

OpenStudy (sriram):

that is if twice the smaller number is 9 less than the larger number

OpenStudy (anonymous):

i assume u do so 2x = y - 9 y - x = 15 adding eliminates y: x = 6 and plug x=6 into second equation y - 6 = 15 y = 21 answer (6,21)